OpenStudy (anonymous):

Two standard dice are rolled. Let A be the event that the first die lands on a number LESS than 4. Let B be the event that the second die lands on 5. Select all that apply. A. P(A and B) = 1/12 B. P(A and B) = 1/36 C. P(A) = 1/2 D. P(B) = 1/6

OpenStudy (anonymous):

I kind of have a hunch that B is not one

OpenStudy (anonymous):

I don't see where the 36 comes from

OpenStudy (anonymous):

you're right. they are actually all right except B

OpenStudy (anonymous):

P(A) = 3/6 = 1/2 P(B) = 1/6 P(A and B) = 1/2 * 1/6 = 1/12

OpenStudy (anonymous):

Ohh wow that's exactly what I was thinking. FINALLY I'm right lol